# Quantifying the Limits of Segmentation Foundation Models: Modeling Challenges in Segmenting Tree-Like and Low-Contrast Objects

### v2

@@ Abstract
Image segmentation foundation models (SFMs) like Segment Anything Model (SAM) have achieved impressive zero-shot and interactive segmentation across diverse domains. However, they struggle to segment objects with certain structures, particularly those with dense, tree-like morphology and low textural contrast from their surroundings. These failure modes are crucial for understanding the limitations of SFMs in real-world applications. To systematically study this issue, we introduce interpretable metrics quantifying object tree-likeness and textural separability. On carefully controlled synthetic experiments and real-world datasets, we show that SFM performance (e.g., SAM, SAM 2, HQ-SAM) noticeably correlates with these factors. We link these failures to “textural confusion”, where models misinterpret local structure as global texture, causing over-segmentation or difficulty distinguishing objects from similar backgrounds. Notably, targeted fine-tuning fails to resolve this issue, indicating a fundamental limitation. Our study provides the first quantitative framework for modeling the behavior of SFMs on challenging structures, offering interpretable insights into their segmentation capabilities.
@@ 1 Introduction
Segment Anything Model (SAM) [26], the most widely-used segmentation foundation model (SFM), has demonstrated promising zero-shot and fine-tuned segmentation ability across diverse domains, including biomedical imaging [31, 21, 10], remote sensing [6, 37], and others. However, certain failure modes for SAM have been found empirically, where SAM produces underwhelming performance on datasets with certain uncommon objects or atypical contexts, such as dense, branching structures (e.g., retinal blood vessels) [38, 11], concealed/low-contrast objects [7, 41, 23, 44], small or irregular objects, and others [23].
@@
We will show that newer SFMs, including SAM 2 [36] and HQ-SAM [24], are also susceptible to these failure modes, despite some stated improvements over these limitations for SAM—HQ-SAM on thin, branching structures and SAM 2 on low-contrast objects. Fine-tuning or adapting these models for such challenging examples using various strategies [30, 33, 16, 24] (or utilizing specialized techniques, e.g., matting algorithms [28]) could mitigate these issues, yet we will show that these strategies do not fully resolve the problems, nor explain the root causes. This underscores the need for a theoretical understanding of why SFMs struggle with such cases, paving the way for better model design and adaptation for new applications.
@@
Figure 1: SAM’s segmentation performance tends to drop noticeably when the object has high tree-likeness (left, on DIS) or low textural separability (right, on iShape)–even with heavy prompting–which we investigate in this work. From left to right: input images, object annotations, and their SAM segmentations.
@@
Here, we propose that several of these failure modes in SFMs stem from misinterpreting textural cues when distinguishing objects from their surroundings. Specifically, SFMs often confuse dense, tree-like structures as large-scale textures, leading to over-segmentation, and struggle with objects that have similar textures to their surroundings. To quantify these challenges, we introduce interpretable metrics for object tree-likeness and textural separability. Tree-likeness (Sec. 4.2) captures structural complexity using spatial contour frequency and the difference in object variability between global and local scales. Textural separability (Sec. 5.1) measures how distinguishable an object’s textural features are from those of its background. These metrics are designed for broad applicability across datasets and object types and are implemented in a PyTorch-compatible form for efficiency.
@@
We then investigate how tree-likeness and textural separability affect SFM performance through controlled synthetic experiments (Secs. 4.3.1, 5.2.1) and real-world datasets (Secs. 4.3.2, 5.2.2). These experiments systematically isolate the influence of object shape and texture while minimizing confounding factors. Across all datasets, we find that SFM performance (IoU) noticeably correlates with both tree-likeness and textural separability across diverse object and image types. Notably, HQ-SAM, designed for intricate structures, still exhibits these limitations, highlighting the persistence of these challenges. Furthermore, our fine-tuning experiments (Sec. 6) confirm that these failure modes remain robust to fine-tuning, even when explicitly trained on the most difficult cases. Our overall contributions are as follows:
@@
    We introduce quantitative metrics for object tree-likeness—Contour Pixel Rate (CPR) and Difference of Gini Impurity Deviation (DoGD)—and an object textural separability metric based on neural feature contrast.
@@
Our findings provide the first quantitative model linking SFM performance to measurable object characteristics. Despite the complexity of SFMs and their evaluation datasets, we show that their underperformance on challenging cases can be modeled using our straightforward, interpretable object descriptors. We release easy-to-use code for our proposed metrics at https://github.com/mazurowski-lab/SAM-TexturalConfusion-Metrics.
@@ 2 Related Work
In addition to the various “failure modes” and biases of segmentation foundation models discussed in the introduction which motivated this work, various works have also studied the same for general vision models. Central to these is the goal of understanding the visual features that networks use to make predictions from images, and whether object shape or texture is more important. This is important to study because any unexpected failure of such models will depend on the features used for inference.
Zhang et al. [49] posited that objects within images can be described by three characteristics: shape, texture, and the composition of textures within shapes, and analyzed how these characteristics may affect the challenge of segmenting such objects. It is well known that convolutional neural networks’ (CNNs’) predictions are typically biased towards the texture of objects rather than their shape, for both classification [15, 40, 20, 2] and segmentation [49], resulting in curious examples where network predictions are based on texture even when the shape clearly defines a different object. For example, [27] found that when a CNN is presented with images of objects with only their shape visible, not texture (silhouettes), performance was greatly worsened.
It has also been found that the shape-learning behavior of neural networks can differ noticeably depending on the particular training dataset and objective [20, 19], motivating us to study the specific behavior of SAM (given its unique training set) and other similar SFMs in detail. Interestingly, there has been some evidence that vision transformers (like SAM’s image encoder) may instead be biased towards shape, rather than texture for classification tasks [42]. Here, we study segmentation rather than classification, where it is unclear if those results extend, as our results generally point to SAM focusing on texture over shape (to the point of even confusing complex, locally varying shape itself as texture), which aligns with the preliminary findings of [48].
@@ 3 Methods: SFM Usage and Prompting
We experiment with the ViT-H (default, if unstated) and ViT-B standard pre-trained SAM models [26], the ViT-L and ViT-B+ variants of SAM 2 [36], and the ViT-H and ViT-B variants of HQ-SAM [24], a SAM adaptation designed particularly for thin structures. As is the default for SAM, all input images are resized to a resolution of $1024\times 1024$, and normalized to $[0,255]$. All images that possess instance segmentations for multiple objects will be evaluated by the given SFM segmenting each of the images’ objects one at a time (guided by clear prompting described in the next paragraph), with all other mask pixels set as background. In all experiments, we use the oracle-chosen mask out of the given SFM’s three outputted predictions.
@@ 4.1 Motivation
We first wished to understand the challenges for SFMs in segmenting certain objects with dense, tree-like structures, and then quantify the features driving these challenges, following our observation of a trend of SAM having difficulty with such objects, including retinal blood vessels [35, 30] or satellite road images [13, 45] (see e.g., Fig. 2). Interestingly, while both of these object types have branching features, a characteristic which naively could relate to the SFM’s “failure”, SAM’s performance on retinal vessel images is noticeably worse (avg. $\mathrm{IoU}\simeq 0.05$ [35]) compared to on satellite road structures (avg. $\mathrm{IoU}\simeq 0.2$ [13]), which was reproduced in our own experiments (Fig. 4) [^1] This could potentially simply be due to retinal vessels being thinner than road structures, but our experiments do not support this (App. D.2)..
@@
We hypothesize that such objects become more challenging to segment when their branching structures are dense and irregularly-spaced, which we refer to as “tree-like”. We will show that SFMs consider such structures as textures rather than shapes, resulting in significant over-segmentation. In the following section, we propose how to characterize the tree-likeness of objects using two new quantitative metrics: CPR (Contour Pixel Rate) and DoGD (Difference of Gini Impurity Deviation).
@@ 4.2.1 Contour Pixel Rate
Definition 1 (Contour Pixels).
Given some mask $m$, a pixel $m_{ij}$ is a contour pixel if $m_{ij}=1$ and there exists a different pixel $m_{kl}$ such that $m_{kl}\neq m_{ij}$ and $||(k,l)-(i,j)||_{1}<R$, given some small contour width threshold $R>0$.
@@
Intuitively, tree-like objects will have a higher percentage of contour pixels, resulting in higher CPRs. We demonstrate fast computation of CPR in full detail in Algorithm 1, via vectorized PyTorch-like pseudocode.
@@ 4.2.2 Difference of Gini-impurity Deviation (DoGD)
$1-[p(\mathcal{W}^{k}_{h_{0},w_{0}}(m))]^{2}-[1-p(\mathcal{W}^{k}_{h_{0},w_{0}}(m))]^{2},$
[…]
@@
Intuitively, objects with significant tree-like or fractal-like structure will exhibit relatively large values of $\sigma^{\mathrm{Gini}}_{b}(m)$ due to high variability in pixel composition at small scales (frequently alternating between areas of mixed classes and areas of a single class), yet small $\sigma^{\mathrm{Gini}}_{a}(m)$ due to structures with high uniformity and/or repetitions at large scales, altogether increasing the DoGD.
We perform all experiments with the hyperparameters for CPR and DoGD set to $R=5$, $a=127$, and $b=3$, which we found via grid search for the values that resulted in the Kendall’s $\tau$ between IoU and DoGD with the lowest $p$-value on the held-out DIS training set using ViT-H SAM. We show results using a wide range of other values for these hyperparameters in Appendix B.1, where our findings were consistent for most other settings. Moreover, we note that CPR and DoGD are correlated with each other (Pearson $|r|=0.83$ on average; Appendix D.1), showing their consistency in quantifying different aspects of tree-likeness.
@@ 4.3.1 Experiments on Synthetic Data
In this section, we will first probe the effect of object tree-likeness on the performance of SFMs by testing them on synthetic images that solely possess objects of varying tree-likeness, with different independently chosen, uniform foreground and background textures. These objects are contiguous components sampled from retinal blood vessel and satellite road masks; the full procedure of generating these images is detailed in Appendix A.2. Example generated images, masks, and SAM predictions for them are shown in Fig. 3. As shown (also in Fig. 4), the objects cover a wide range of tree-likeness as measured by these quantities.
@@
In Fig. 4 we show the relationship between an object’s tree-likeness (via CPR or DoGD) with SAM’s performance (IoU) in segmenting the object, for all generated synthetic objects. We quantify the strength of this relation via non-linear correlation, measured by Kendall’s tau ($\tau$) [25] and Spearman’s rho ($\rho$) [39], for all three SFMs, shown in Table 1. We see that object tree-likeness is quite predictive of an SFM’s performance, with average absolute correlations of $|\tau|=0.75$ and $|\rho|=0.92$ for CPR and $|\tau|=0.67$ and $|\rho|=0.84$ for DoGD; i.e., more prevalent dense tree-like structures corresponding to worse performance.
@@
Table 1: Rank correlation between SFM IoU and object tree-likeness (CPR and DoGD), on the synthetic dataset.
Figure 4: SAM IoU vs. object tree-likeness (CPR, left; DoGD, right) on the synthetic dataset.
While this relationship is certainly strong, it is still on synthetic, controlled data. In the following section, we will show that this finding is also present for real data which is subject to noise from a variety of uncontrollable factors.
@@ 4.3.2 Experiments on Real Data
We will now perform the same analysis on two real datasets which contain objects which cover a spectrum of tree-likeness, DIS5k and iShape. DIS5k [34] (or “DIS”) is a dataset containing extremely detailed annotations of objects with varying degrees of hollowness, and both regular and irregular tree-like structures. All DIS experiments will be reported on its validation set unless otherwise stated. We show example images with objects with varying degrees of tree-likeness (by CPR) in Fig. 1 left.
iShape [46] consists of six sub-datasets of real and realistic-appearing synthetic images for instance segmentation of different objects: antenna, branches, fences, logs, hangers, and wires (see e.g. Fig. 13). We analyze these classes individually to mitigate potential confounding/noise factors due to inter-class variations; we do not do this for DIS due to the larger number of classes which are much more fine-grained in their differences, additionally because there are few images per class for DIS.
In Fig. 5 we show how SAM’s performance (IoU) on these images relates to the tree-likeness of the objects which it is segmenting, alongside the other SFMs (SAM 2 and HQ-SAM) for comparison for DIS, with accompanying quantitative correlation results for all three SFMs shown in Table 2. In order to reduce the noise incurred by the large variety of segmented objects in the dataset and nuisance confounding factors in the images, we analyze results with aggregated objects: we cluster groups of $5$ objects/images with similar IoU and tree-likeness (CPR or DoGD) into single datapoints of the average value of these metrics, and similar for the IoU vs. textural separability experiments of Sec. 5.

### v3

@@ Abstract
Image segmentation foundation models (SFMs) like Segment Anything Model (SAM) have achieved impressive zero-shot and interactive segmentation across diverse domains. However, they struggle to segment objects with certain structures, particularly those with dense, tree-like morphology and low textural contrast from their surroundings. These failure modes are crucial for understanding the limitations of SFMs in real-world applications. To systematically study this issue, we introduce interpretable metrics quantifying object tree-likeness and textural separability. On carefully controlled synthetic experiments and real-world datasets, we show that SFM performance (e.g., SAM, SAM 2, HQ-SAM) noticeably correlates with these factors. We attribute these failures to SFMs misinterpreting local structure as global texture, resulting in over-segmentation or difficulty distinguishing objects from similar backgrounds. Notably, targeted fine-tuning fails to resolve this issue, indicating a fundamental limitation. Our study provides the first quantitative framework for modeling the behavior of SFMs on challenging structures, offering interpretable insights into their segmentation capabilities. [^1] Code: https://github.com/mazurowski-lab/SAMFailureMetrics
@@ 1 Introduction
Segment Anything Model (SAM) [sam], the most widely-used segmentation foundation model (SFM), has demonstrated promising zero-shot and fine-tuned segmentation ability across diverse domains, including biomedical imaging [MAZUROWSKI2023102918, huang2024segment, osti_10447851], remote sensing [chen2024rsprompter, ren2024segment], and others. However, certain failure modes for SAM have been found empirically, where SAM produces underwhelming performance on datasets with certain uncommon objects or atypical contexts, such as dense, branching structures (e.g., retinal blood vessels) [shi2023generalist, dong_efficient_2024], concealed/low-contrast objects [chen2023sam, tang2023can, ji2024segment, wang2025orderaware], small or irregular objects, and others [ji2024segment].
@@
We will show that newer SFMs, including SAM 2 [sam2] and HQ-SAM [hqsam], are also susceptible to these failure modes, despite some stated improvements over these limitations for SAM—HQ-SAM on thin, branching structures and SAM 2 on low-contrast objects. Fine-tuning or adapting these models for such challenging examples using various strategies [ma2024segment, peng2024sam, gu2024build, hqsam] (or utilizing specialized techniques, e.g., matting algorithms [lin2021real]) could mitigate these issues, yet we will show that such strategies do not fully resolve the problems, nor explain the root causes. This underscores the need for a theoretical understanding of why SFMs struggle with such cases, paving the way for better model design and adaptation for new applications.
@@
Figure 1: SAM’s segmentation performance tends to drop noticeably when the object has high tree-likeness (left, on DIS) or low textural separability (right, on iShape)–even with heavy prompting–which we investigate in this work.
@@
Here, we propose that several of these failure modes in SFMs stem from misinterpreting textural cues when distinguishing objects from their surroundings. Specifically, SFMs often confuse dense, tree-like structures as large-scale textures, leading to over-segmentation, and struggle with objects that have similar textures to their surroundings. To quantify these challenges, we introduce interpretable metrics for object tree-likeness and textural separability, designed to quantify specific aspects of objects that may make them challenging to segment by typical SFM architectures. Tree-likeness (Sec. 4.2) captures structural complexity with two new metrics, involving spatial contour frequency and the difference in object variability between global and local scales, inspired by how vision transformer encoders and attention patterns in SFMs can be “confused” by tree-like objects. Textural separability (Sec. 5.1) measures how distinguishable an object’s textural features are from those of its background. These metrics are designed for broad applicability across datasets and objects and are implemented in a PyTorch-compatible form for efficiency.
@@
We then investigate how tree-likeness and textural separability affect SFM performance through controlled synthetic experiments (Secs. 4.3.1, 5.2.1) and real-world datasets (Secs. 4.3.2, 5.2.2). These experiments systematically isolate the influence of object shape and texture while minimizing confounding factors. Across all datasets, we find that SFM performance (segmentation IoU) noticeably correlates with both tree-likeness and textural separability across diverse object and image types. Notably, HQ-SAM, designed for intricate structures, still exhibits these limitations, highlighting the persistence of these challenges. Furthermore, our fine-tuning experiments (Sec. 6) confirm that these failure modes remain robust to fine-tuning, even when explicitly trained on the most difficult cases. Our overall contributions are as follows:
@@
    We introduce quantitative novel metrics for object tree-likeness—Contour Pixel Rate (CPR) and Difference of Gini Impurity Deviation (DoGD)—and an object textural separability metric based on neural feature contrast.
@@
Our findings provide the first quantitative modeling that links SFM performance to measurable object characteristics. Despite the complexity of SFMs and their evaluation datasets, we demonstrate that their underperformance on challenging cases can be effectively modeled using our straightforward and interpretable object descriptors.
@@ 2 Related Work
Beyond the original Segment Anything Model (SAM) [sam] and the subsequent SFMs of HQ-SAM [hqsam] and SAM 2 [sam2], other works involving universal, generalist, and/or foundation models or architectures for segmentation include OneFormer [jain2023oneformer], a framework for multi-task universal image segmentation; Segment and Caption Anything (SCA), a method equipping SAM with caption generation [huang2024sca]; Semantic-SAM [li2024semanticsam], for segmenting with controllable granularity between objects; Open-Vocabulary SAM [yuan2024open], a framework unifying SAM and CLIP [clip]; and SEEM [zou2023seem], a less commonly-used multimodal foundation model developed concurrently with SAM.
Previous work on biases and failure modes in computer vision models for tasks such as segmentation has primarily focused on understanding whether networks rely more on object shape or texture for making visual predictions. Zhang et al. [zhang2024convolutional] proposed a framework describing objects by shape, texture, and texture composition within shapes, highlighting how these characteristics impact segmentation tasks. Extensive literature shows that convolutional neural networks (CNNs) tend to be biased toward texture over shape, affecting performance in both classification [geirhosimagenet2019, subramanian2024spatial, hermann2020origins, baker2018deep] and segmentation [zhang2024convolutional] tasks. Notably, CNN performance significantly drops when tested on shape-only images (silhouettes), underscoring texture bias [kubilius2016deep].
Moreover, neural networks’ shape-learning behaviors vary based on the training dataset and objectives [hermann2020origins, hermann2020shapes]. This variation motivates our detailed investigation into the specific behaviors of SFMs such as the Segment Anything Model (SAM), given their unique training characteristics. While vision transformers (such as SAM’s image encoder architecture) have demonstrated a preference for shape in classification tasks [tuli2021convolutional], it remains uncertain if this extends to segmentation. Our results suggest SAM predominantly leverages texture cues, sometimes even interpreting complex shapes as textures, corroborating prior studies [zhang2023understanding].
@@ 3 Methods: SFM Usage and Prompting
We experiment with the ViT-H (default, if unstated) and ViT-B standard pre-trained SAM models [sam] (labeled “SAM-H” and “SAM-B”), the ViT-L and ViT-B+ variants of SAM 2 [sam2] (“SAM 2-L” and “SAM 2-B+”), and the ViT-H and ViT-B variants of HQ-SAM [hqsam] (“HQ-SAM-H” and “HQ-SAM-B“), a SAM adaptation designed particularly for thin structures. As is the default for SAM, all input images are resized to a resolution of $1024\times 1024$, and normalized to $[0,255]$. All images that possess instance segmentations for multiple objects will be evaluated by the given SFM segmenting each of the images’ objects one at a time (guided by clear prompting described in the next paragraph), with all other mask pixels set as background. In all experiments, we use the oracle-chosen prediction from the given SFM.
@@ 4.1 Motivation
We first wished to understand the challenges for SFMs in segmenting certain objects with dense, tree-like structures, and then quantify the features driving these challenges, following our observation of a trend of SAM having difficulty with such objects, including retinal blood vessels [qiu2023learnable, ma2024segment] or satellite road images [feng2024road, xu2023leveraging] (see e.g., Fig. 2). Interestingly, while both of these object types have branching features, a characteristic which naively could relate to the SFM’s “failure”, SAM’s performance on retinal vessel images is noticeably worse (avg. $\mathrm{IoU}\simeq 0.05$ [qiu2023learnable]) compared to on satellite road structures (avg. $\mathrm{IoU}\simeq 0.2$ [feng2024road]), which was reproduced in our own experiments (Fig. 4) [^2] This could potentially simply be due to retinal vessels being thinner than road structures, but our experiments do not support this (App. D.4)..
@@
We hypothesize that objects with dense and irregularly-spaced branching structures–which we refer to as ”tree-like”–pose fundamental challenges to SFMs because their structural properties create a mismatch with patch-based attention mechanisms. When a ViT [dosovitskiy2020image] divides an object into fixed-size patches, tree-like structures present a critical issue: their boundary pixels are distributed throughout the object rather than concentrated at the perimeter. Consequently, most patches contain predominantly boundary information, leaving insufficient object-internal features to build coherent representations. In the following section, we will formalize this via our proposed Contour Pixel Rate (CPR) metric. SFM self-attention maps for tree-like objects for SAM then have more fragmented, locally-focused attention patterns that respond to local edges independently rather than forming coherent global object representations, which we show empirically (via the spatial autocorrelation of attention maps) in App. D.2.2 with examples shown in App. D.1. Moreover, we empirically show that typical object CPR increases with a model’s effective patch size (App. D.2.1).
Beyond boundary density, tree-like structures also exhibit dramatic appearance variation across different scales: thick trunks, medium branches, and fine tendrils create inconsistent local patterns. We capture this through our Distance of Gini impurity Deviation (DoGD) metric, proposed in the following section, which measures how object pixel distribution varies across scales. High DoGD values indicate scale-dependent spatial inconsistency, preventing stable object representations.
The combination of high CPR and high DoGD explains why SFMs fail on tree-like structures: boundary-dominated patches (high CPR) produce fragmented attention lacking semantic coherence, while scale-dependent variations (high DoGD) prevent consistent object-level representations. This causes SFMs to interpret tree-like objects as textures–collections of local edges without global structure–leading to over-segmentation where unified objects are split into multiple regions. Our metrics thus provide a quantitative framework for predicting SFM failures on complex branching structures.
@@ 4.2.1 Contour Pixel Rate
Definition 4.1 (Contour Pixels).
Given some mask $m$, a pixel $m_{ij}$ is a contour pixel if $m_{ij}=1$ and there exists a different pixel $m_{kl}$ such that $m_{kl}\neq m_{ij}$ and $(k,l)-(i,j)||_{1}<R$, given some smallthreshold $R>0$.
@@
The more tree-like an object is, the higher the percentage of its pixels that are contour pixels, resulting in a higher CPR. We demonstrate fast computation of CPR in full detail in Algorithm 1, via vectorized PyTorch-like pseudocode.
@@ 4.2.2 Difference of Gini-impurity Deviation (DoGD)
$1-[p(\mathcal{W}^{k}_{h_{0},w_{0}}(m))]^{2}-[1-p(\mathcal{W}^{k}_{h_{0},w_{0}}(m))]^{2}$
where we write $p:=p_{1}$. The Gini impurity measures the degree of uncertainty (ranging from 0 to $0.5$) of whether an object is present in the given window. For example, mask windows containing pixels of mostly one class will have $\mathrm{Gini}\simeq 0$, while having similar pixel amounts of both classes will result in $\mathrm{Gini}\simeq 0.5$ [^3] Following common decision tree practices, we use Gini impurity over entropy because it is a symmetric and faster alternative [breiman1984classification]..
@@
Objects with significant tree-like or fractal-like structure will exhibit relatively large values of $\sigma^{\mathrm{Gini}}_{b}(m)$ due to high variability in pixel composition at small scales (frequently alternating between areas of mixed classes and areas of a single class), yet small $\sigma^{\mathrm{Gini}}_{a}(m)$ due to structures with high uniformity and/or repetitions at large scales, altogether increasing the DoGD.
We perform all experiments with the hyperparameters for CPR and DoGD set to $R=5$, $a=127$, and $b=3$, which we found via grid search for the values that resulted in the Kendall’s $\tau$ between IoU and DoGD with the lowest $p$-value on the held-out DIS training set using SAM-H. We show results using a wide range of other values for these hyperparameters in App. B.1, where our findings were consistent for most other settings. Moreover, we note that CPR and DoGD are noticeably correlated (Pearson $|r|=0.83$ on average; App. D.3), showing their consistency in quantifying different aspects of tree-likeness.
@@ 4.3.1 Experiments on Synthetic Data
In this section, we will first probe the effect of object tree-likeness on the performance of SFMs by testing them on synthetic images that solely possess objects of varying tree-likeness, with different independently chosen, uniform foreground and background textures. These objects are contiguous components sampled from retinal blood vessel and satellite road masks; the full procedure of generating these images is detailed in App. A.2. Example generated images, masks, and SAM predictions for them are shown in Fig. 3. As shown (also in Fig. 4, left), the objects cover a wide range of tree-likeness as measured by these quantities.
@@
In Fig. 4 left, we show the relationship between an object’s tree-likeness (via CPR or DoGD) with SAM’s performance (IoU) in segmenting the object, for all generated synthetic objects. We quantify the strength of this relation via non-linear correlation, measured by Kendall’s tau ($\tau$) [kendall1938new] and Spearman’s rho ($\rho$) [spearman_proof_1904], for all three SFMs, shown in the table to the right of the plots. We see that object tree-likeness is quite predictive of an SFM’s performance, with average absolute correlations of $|\tau|=0.75$ and $|\rho|=0.92$ for CPR and $|\tau|=0.67$ and $|\rho|=0.84$ for DoGD; i.e., more prevalent dense tree-like structures corresponding to worse performance. In the following section, we will show that this finding is also present for real data, which is subject to noise from a variety of uncontrollable factors.
Refer to captionRefer to caption
@@
Figure 4: Left: SFM prediction IoU vs. object tree-likeness (CPR and DoGD), on the synthetic dataset, shown for SAM-H. Right: Rank correlations between IoU and tree-likeness for all SFMs.
@@ 4.3.2 Experiments on Real Data
We will now perform the same analysis on three real datasets which contain objects which cover a spectrum of tree-likeness, DIS5k, MOSE, and iShape. DIS5k [DIS] (or “DIS”) is a dataset containing extremely detailed annotations of objects with varying degrees of hollowness, and both regular and irregular tree-like structures. All DIS experiments will be reported on its validation set. We show example images with objects with varying degrees of tree-likeness (by CPR) in Fig. 1 left. We additionally performed experiments on the MOSE dataset [MOSE] of videos of objects in complex, cluttered, and/or camouflaged scenes, randomly sampled 4 frames with non-empty labels from each of the 1,507 videos in the training set, resulting in $\sim$6,000 frames with $\sim$13,000 total objects.
iShape [yang2021ishape] consists of six sub-datasets of real and realistic-appearing synthetic images for instance segmentation of different objects: antenna, branches, fences, logs, hangers, and wires (see e.g. Fig. 13). We analyze these classes individually to mitigate potential confounding/noise factors due to inter-class variations.
In Fig. 5 left, we show how SAM’s performance (IoU) on these images relates to the tree-likeness of the objects which it is segmenting, alongside the other SFMs (SAM 2 and HQ-SAM) for comparison for DIS, with accompanying quantitative correlation results for all three SFMs shown in the table to the right of the plots. The plots for MOSE results are provided in App. D.7 due to space constraints. In order to reduce the noise incurred by the large variety of segmented objects in the dataset and nuisance confounding factors in the images, we analyze results with aggregated objects: we cluster groups of $5$ objects/images ($20$ for MOSE due to its size) with similar IoU and tree-likeness (CPR or DoGD) into single datapoints of the average value of these metrics, and similar for the IoU vs. textural separability experiments of Sec. 5.
Refer to captionRefer to captionRefer to captionRefer to caption